Circular No. 7511

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2000dw IN UGC 11955 Y. L. Qiu and J. Y. Hu, Beijing Astronomical Observatory (BAO), on

behalf of the BAO Supernova Survey, report the discovery of an apparent supernova (mag ∼ 17.5) in unfiltered images taken on Oct. 17.5 and 20.6 UT with the BAO 0.6-m telescope. The new object was confirmed by W. D. Li, University of California at Berkeley, on unfiltered images taken on Oct. 19.2 with the 0.8-m Katzman Automatic Imaging Telescope. SN 2000dw is located at α = 22^{h}13^{m}49^{s}.06, δ = +39^{o}14′15′′.1 (equinox 2000.0), which is 1′′.5 east and 3′′.3 north of the center of UGC 11955. Photometry is difficult, as the object is very close to the nucleus of the parent galaxy. Unfiltered BAO images taken on July 12 (limiting mag ∼ 19.0) showed no star at the position of the new object.

SUPERNOVAE 2000ds, 2000dt, 2000du, 2000dv A. V. Filippenko and R. Chornock, University of California at Berkeley,

communicate: “Inspection of CCD spectra (range 330–1000 nm) of SN 2000dt (IAUC 7508), obtained on Oct. 24 UT with the Shane 3-m reflector at Lick Observatory, reveal that the object is probably of type Ib. He I absorption lines are present, but there is also an unidentified line at rest wavelength 622 nm (possibly blueshifted C II 658 nm, or perhaps Hα, though the corresponding Hβ line seems absent). SN 2000dv (IAUC 7509) is also a type-Ib supernova, ∼ 3 months past maximum brightness; it is starting to enter the nebular phase, but He I absorption is still present. SN 2000ds (IAUC 7507) is an even older type-Ib (or -Ic) supernova, perhaps 4 to 7 months past maximum; we prefer the type-Ib classification because the [Ca II] 732-nm emission has FWHM ∼ 5000 km/s, most consistent with type-Ib supernovae (Filippenko et al. 1995, Ap.J. 450, L11; Matheson et al. 2001, A.J., in press). SN 2000du (IAUC 7509) is a normal type-II supernova; Hα exhibits a classic P-Cyg profile.”

COMET C/1999 Y1 (LINEAR) Total visual magnitude estimates (L = reflector): Aug. 5.00 UT, 13.7

(M. Meyer, Frauenstein, Germany, 0.25-m L); 26.30, 14.2 (A. Hale, Cloudcroft, NM, 0.41-m L); Sept. 8.90, 13.7 (W. Hasubick, Buchloe, Germany, 0.44-m L); Oct. 6.01, 12.9 (G. Comello, Roden, The Netherlands, 0.31-m L); 17.79, 12.6 (R. J. Bouma, Groningen, The Netherlands, 0.25-m L); 23.76, 12.1 (M. Lehký, Hradec Králové, Czech Rep., 0.42-m L).

2000 October 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green